HUBBLE SPACE TELESCOPE DAILY REPORT #2469 PERIOD COVERED: 0000Z (UTC) 09/30/99 - 0000Z (UTC) 10/01/99 Daily Status Report as of 274/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.4 Completed STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.5 Completed WF/PC-2 8441 (WF/PC-2 Decontaminations and Associated Observations Pt. 1/2) The WF/PC-2 was used to make its monthly decontamination operations. Also included are instrument monitors tied to decons: photometric stability check, focus monitor, pre- and post-decon internals {bias, internal flats, k-spots, and darks}, and ultraviolet throughput checks. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 8313 (Imaging the Host Galaxies of High Redshift Type Ia Supernovae) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of 972 and its host galaxy. The propose is to obtain snapshot, unfiltered STIS images of distant galaxies of known redshift which have hosted supernovae {SNe} of Type IA found via the Supernova Cosmology Project. The observations were completed as scheduled, and no anomalies were reported. 1.7 Completed STIS/CCD 8065 (Comparison of Protoplanetary Disks) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the protoplanetary disk around star HD163296. Multiple roll angles were used to minimize systematic errors and image artifacts. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed WF/PC-2 8160 (The Optical Counterpart of the Soft Gamma-Ray Repeater 1900+14) The WF/PC-2 was to be used to make observations of NGC-104 to search for the optical counterpart of the soft gamma-ray repeater 1900+14. As documented in 2.1 and HSTAR 7287, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.9 Completed S/C 7725 (STIS NUV MAMA Anomalous Recovery) This proposal is designed to permit recovery of the NUV MAMA detector after an anomalous shutdown. As documented in 2.1 and HSTAR 7287, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.10 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was to be used to make a ultraviolet Earth flat calibration image. As documented in 2.1 and HSTAR 7287, the acquisition/re-acquisitions for both iterations of this proposal failed and the take data flag remained down. 1.11 Completed STIS/MA2 7643 (STIS MAMA Fold Analysis for the Anomalous Recovery Procedure) The Space Telescope Imaging Spectrograph (MA2) was to be used to conduct fold analysis for the anomalous recovery procedure of the NUV and FUV MAMA detectors. As documented in 2.1 and HSTAR 7287, the acquisition/re-acquisitions for this proposal failed and the take data flag remained down. 1.12 Completed STIS/CCD 8271 (A Snapshot Survey of LMC Planetary Nebulae: A Study of Nebular and Late Stellar Evolution) The Space Telescope Imaging Spectrograph (CCD) was used to make medium-resolution slitless spectroscopy snapshot images of the planetary nebula LMC-SMP-80 located in the Large Magellanic Cloud to study the parallel evolution of the nebulae and their central stars. The observations were completed as planned, and no problems were reported. 1.13 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 8396 (WF/PC-2 Observations of Astrophysically Important Visual Binaries - Continued) The WF/PC-2 was used to make observations of the visible binary system G107-70. These observations promise to add two accurate masses to the tiny handful of white-dwarf masses that are directly known from dynamical measurements. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 9 Successful: 8 Scheduled Re-acquisitions: 3 Successful: 1 Per HSTAR 7287, the acquisition at 273/170523Z failed to RGA control after multiple attempts. The search radius and scan step flags were set for FGS-1. The re-acquisitions at 273/183532Z and at 273/201209Z also failed. The proposals described in 1.8, 1.9, 1.10, and 1.11 were affected. 2.2 FHST Updates: Scheduled: 28 Successful: 27 Per HSTAR 7285, the Roll Delay Update at 273/101741Z failed. 2.3 Operations Notes: Using ROP SR-1, the SSR EDAC error counter was cleared four times. At 273/111453Z and at 273/111523Z, batteries 4 and 2 voltages flagged our-of-limits low, remaining there for two samples at a value of 27.636 volts. The limit is 27.64 volts. HSTAR 7286 was written. The STIS MAMA-2 recovery exercise completed successfully. The SSA transmitter was managed per ROP !C-2, and SSR and ESTR-2 dumps were made. The results of those dumps are being analyzed. Per HSTAR 7288, the battery 2 current flagged out-of-limits high at 274/080245Z with a value of 18.6 amps. The high limit is 18.0 amps. There were type 2 and solar array slews taking place at the time. The load bus current also flagged at this time (97.2 amps vs. the 96.0 amp limit). The NICMOS FPA temperature sensor for detector 3 was enabled at 274/0311Z. Using ROP NS-3, the NSSC-1 status buffer was dumped at 274/0122Z. The CCS “A” string SAC server was re-cycled at 274/0754Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH